### v2

# Precovery Observations of 3I/ATLAS from TESS Suggests Possible Distant Activity

Abstract
3I/ATLAS is the third macroscopic interstellar object detected traversing the Solar System. Since its initial discovery on UT 01 July 2025, hundreds of hours on a range of observational facilities have been dedicated to measure the physical properties of this object. These observations have provided astrometry to refine the orbital solution, photometry to measure the color, a rotation period and secular light curve, and spectroscopy to characterize the composition of the coma. Here, we report precovery photometry of 3I/ATLAS as observed with NASA’s Transiting Exoplanet Survey Satellite (TESS). 3I/ATLAS was observed nearly continuously by TESS from UT 07 May 2025 to 02 June 2025. We use the shift-stack method to create deep stack images to recover the object. These composite images reveal that 3I/ATLAS has an average TESS magnitude of and an absolute visual magnitude of , the latter being consistent with magnitudes reported in July 2025. When coupled with recent HST images deriving a nucleus size of R2.8 km (H15.4), our measurements suggest that 3I/ATLAS may have been active out at au. Additionally, we extract a day light curve and find no statistically significant evidence of a nucleus rotation period. Nevertheless, the data presented here are some of the earliest precovery images of 3I/ATLAS and may be used in conjunction with future observations to constrain the properties of our third interstellar interloper.
show]adina@msu.edu
1 Introduction
Interstellar objects (ISOs) offer unique insights into the formation conditions of rocks and ices in regions of our galaxy beyond the Solar System. The recent discovery of the third ISO, 3I/ATLAS, was reported on UT 01 July 2025 (L. Denneau et al., 2025) by the ATLAS survey (J. L. Tonry et al., 2018a, b). This discovery increases the current census of discovered macroscopic ISOs by 50%. Since its discovery, there has been an international rush to observe 3I/ATLAS before it is no longer visible to observers in February 2026. In particular, early telegrams reported activity in images from the Nordic-Optical Telescope (D. Jewitt & J. Luu, 2025) and the Two-meter Twin Telescope (M. R. Alarcon et al., 2025). Preliminary reconnaissance observations revealed a reddened reflectance spectrum similar to D-type asteroids, no clear photometric variability and faint cometary activity (T. Kareta et al., 2025; R. de la Fuente Marcos et al., 2025; C. Opitom et al., 2025; D. Z. Seligman et al., 2025; B. Yang et al., 2025).
Two macroscopic scale ISOs had been discovered before 3I/ATLAS: 1I/‘Oumuamua in 2017 (G. V. Williams et al., 2017) and 2I/Borisov in 2019 (G. Borisov et al., 2019). The light curve of 1I/‘Oumuamua was markedly different from the 3I/ATLAS because it exhibited drastic brightness variations (M. Drahus et al., 2017; W. C. Fraser et al., 2018; M. M. Knight et al., 2017; M. J. S. Belton et al., 2018). This was presumably due to an elongated oblate spheroid geometry (S. Mashchenko, 2019; A. G. Taylor et al., 2023) and a lack of cometary activity (D. Jewitt et al., 2017; K. J. Meech et al., 2017; Q.-Z. Ye et al., 2017). A significant non-detection of 1I/‘Oumuamua with the Spitzer space telescope provided upper limits on the production of carbon-based species (D. E. Trilling et al., 2018). However, 1I/‘Oumuamua had significant radial nongravitational acceleration in its trajectory (M. Micheli et al., 2018). 1I/‘Oumuamua had a reddened color like 3I/ATLAS (M. T. Bannister et al., 2017; A. Fitzsimmons et al., 2018; J. Masiero, 2017) and a low velocity at infinity indicating a young kinematic age (E. Gaidos et al., 2017; E. Mamajek, 2017). On the contrary, 2I/Borisov displayed distinct cometary activity (D. Jewitt & J. Luu, 2019; A. Fitzsimmons et al., 2019; T. Kareta et al., 2020; C. Opitom et al., 2019; M. T. Bannister et al., 2020; D. Bodewits et al., 2020; G. Cremonese et al., 2020; M. A. Cordiner et al., 2020; P. Guzik et al., 2020; M.-T. Hui et al., 2020; Y. Kim et al., 2020; H. W. Lin et al., 2020; A. J. McKay et al., 2020; Z. Xing et al., 2020; K. Aravind et al., 2021; S. Bagnulo et al., 2021; B. Yang et al., 2021; S. E. Deam et al., 2025). For reviews on ISOs discovered passing through the solar system prior to 3I/ATLAS, we refer the reader to A. Moro-Mart´ın (2022); D. Jewitt & D. Z. Seligman (2023); D. Z. Seligman & A. Moro-Mart´ın (2023); A. Fitzsimmons et al. (2024).
Precovery observations of small bodies such as comets, asteroids, and ISOs can provide constraints on critical physical properties such as orbit, size, composition, and activity. For example, M.-T. Hui & M. M. Knight (2019) searched for precovery detections of 1I/‘Oumuamua in Solar and Heliospheric Observatory (SOHO) and Solar TErrestrial RElations Observatory (STEREO) images. Their search provided nondetections, but these nondetections provided critical upper limits on the production rate of dust and water at low heliocentric distances — during a regime of 1I/‘Oumuamua’s trajectory which was entirely unconstrained from extant observations. Q. Ye et al. (2020) reported pre-discovery observations of 2I/Borisov in survey data from the Zwicky Transient Facility (ZTF), as well as a comprehensive search of Catalina Sky Survey and Pan-STARRS data for the object. The combination of precovery detections and nondetections provided valuable upper limits on the size of the nucleus and information regarding the volatility of ices driving activity. Thus far for 3I/ATLAS, D. Z. Seligman et al. (2025) reported ZTF observations dating back to UT 2025 May 22 and C. O. Chandler et al. (2025) reported observations of using the NSF-DOE Vera C. Rubin Observatory Science Verification (SC) images dating back to UT 2025 June 21. The Rubin Observatory SC observations also revealed faint activity.
All four aforementioned studies highlight the strengths of all-sky surveys in searching for precovery observations. Although not its original science purpose, NASA’s Transiting Exoplanet Survey Satellite (TESS; G. R. Ricker et al., 2015) has also been able to recover and discover small bodies in the solar system (A. Pál et al., 2020; G. M. Szabó et al., 2022; C. Kiss et al., 2025) as well as characterize cometary activity (T. L. Farnham et al., 2019, 2021). TESS has a large on-sky coverage due to the orientation of four wide-field cameras, where each camera has a field-of-view of , amounting to a total FOV of . These images, known as the TESS Full-Frame Images (FFIs), are made available to the community. TESS observes a single field for days, providing a long photometric baseline for variability studies. While limited in its resolution for morphology analysis of cometary comae due to its large pixel scale, TESS is excellent for long-term photometric monitoring and analysis.
[…]
2 TESS Observations
TESS re-observed part of the ecliptic plane during Cycle 7 of its second extended mission. Upon discovery of 3I/ATLAS, we used tess-point (C. J. Burke et al., 2020) in conjunction with 3I/ATLAS’ best-fit orbital solution from the JPL Horizons Small Bodies Database111https://ssd.jpl.nasa.gov/ to determine if the object was in the TESS FOV. We found that 3I/ATLAS was observed during Sector 92, which occurred from UT 07 May through 03 June 2025.
Each of TESS’s four cameras are divided into four charged-coupled devices (CCDs). Due to its high velocity, 3I/ATLAS was observed on two Camera-CCD pairings throughout the sector: Camera 2 CCD 3 and Camera 1 CCD 2. These are located next to each other in the orientation presented in Figure 1. Each TESS FFI has an exposure time of seconds. The data is downloaded every days from the spacecraft due to the increased data volume of the higher cadence FFIs. These downlink times occur near apogee and perigee of the orbit of the spacecraft and thereby result in short gaps in the data. The total duration of science data collected during Sector 92 is 26.05 days.
2.1 FFI Selection and Background Removal
We downloaded all calibrated FFIs for Camera 2 CCD 3 and Camera 1 CCD 2 using the bulk download scripts provided by the Mikulski Archive for Space Telescopes (MAST)222 https://archive.stsci.edu/missions-and-data/tess. 3I/ATLAS was observed on pixels on Camera 2 CCD 3 and pixels on Camera 1 CCD 2. We highlight example images of these two regions in Figure 1. It is important to note that columns are serial register columns, columns are virtual columns, and rows are buffer, smear, and virtual rows; none of these regions can be used for science.
The TESS FFIs are contaminated by both zodiacal light and scattered earthshine contamination. In particular, the scattered earthshine manifests as strong ramps at the beginning and end of each TESS orbit ( days). While the background typically varies smoothly, there have been instances of more sporadic and variable behavior (Figure 2). To identify highly contaminated FFIs, we calculated the median value in each FFI . The FFIs containing highly non-smooth variability, as identified by-eye, are entirely removed from our analysis (Figure 2). This resulted in the removal of 9% of the available FFIs from Camera 2 CCD and 5% of the available FFIs from Camera 1 CCD 2. In total, after these data-quality cuts, we analyzed 9837 FFIs (353 GB). We provide more specific details pertaining to each Camera/CCD configuration in Table 1.
We fit and subtract the background from all remaining FFIs after the heavily contaminated ones were removed. Working with the entire FFI is unnecessary. Instead, we create a “postcard” cutout (A. D. Feinstein et al., 2019) of the entire region where 3I/ATLAS is visible in the FFI (orange boxes in the upper two panels and full bottom two panels in Figure 1). Previous search and recovery analyses of known solar system bodies with TESS have fit pixel variability using an nth-order polynomial (e.g. C. Kiss et al., 2025). We experimented with two methods to remove background contamination from these observations. First, we fit a 2nd-order polynomial to the background, similar to C. Kiss et al. (2025). We fit each continuous section of the light curve with these models, i.e. we chunk the observations based on gaps in the TESS data. This leaves a total of five different segments that are fit (two segments for Camera 2 CCD 3; three segments for Camera 1 CCD 2). This experiment demonstrated that a 2nd-order polynomial fits the baseline well, but fails to fit the orbital ramps (Figure 10).
Second, we fit the background with a data-driven Savitsky-Golay filter with a pre-defined window length and a 2nd-order polynomial. We repeat the same process of chunking the light curve based on gaps in the observations. We test several window lengths ranging from 31 - 901 pixels. The residuals of several example model fits are presented in Figure 10. This experiment revealed that smaller window sizes removed astrophysical signals, while larger windows did not capture the variability of the orbital ramps. Therefore, we settled on using an average window length of 307 pixels.
| Camera/ CCD | 2/3 | 1/2 |
| Start [TBJD] | 3802.975926 | 3815.799580 |
| End [TBJD] | 3813.474497 | 3828.911416 |
| Start [UTC] | 2025-05-07 11:24:43 | 2025-05-20 07:10:47 |
| End [UTC] | 2025-05-17 23:22:39 | 2025-06-02 09:51:49 |
| x pixel range | [1592, 1635] | [556, 601] |
| y pixel range | [0, 305] | [39, 614] |
| [au] | 6.35 - 5.99 | 5.92 - 5.47 |
| [au] | 6.36 - 6.02 | 5.92 - 5.48 |
| 4399 | 5438 | |
2.2 Removing Crowded Frames
Due to the crowdedness in some of the images, we create a filter to remove images which may have significant stellar contamination. The final deepstack images in this work only include frames that are not significantly contaminated. The contamination criteria is defined as the following: Because we are only interested in crowding near the location of 3I/ATLAS, we only consider sources within an window around the central location. We use astropy.stats.sigma_clip to determine how many pixels are outliers compared to the mean value. We assume the default settings for astropy.stats.sigma_clip, other than defining the value. We mask any frame for which more than half of the pixels within this region are flagged as outliers. Examples of non-crowded versus crowded frames are shown in Figure 3.
2.3 Deep Image Creation
After the background models are removed, we employ a shift-stack algorithm to create a TESS deepstack image. We choose not to combine images across detectors, as they may have different systematics. Our algorithm works as follows: (I) Query the location of 3I/ATLAS from the JPL Horizons Small Bodies Database for each FFI; (II) Use the world coordinate system information to convert the queried (RA, Dec) to the (x,y) pixel location of 3I/ATLAS. We convert the (x, y) location from float values to integers using np.round, in order to accurately determine the nearest pixel; (III) Create a pixel cutout centered on the integer (x, y) coordinates where 3I/ATLAS is expected to be; (IV) Repeat this procedure for all integrations; (V) Sum all of the images. Additionally, due to the apparent location with respect to the galactic plane, we create a data quality flag to track particularly crowded fields. The shift-stack technique implemented in this paper works well for objects with known orbital parameters, but may be more computationally inefficient when completing a blind-search for new objects. However, it has been employed with some success for example by W. C. Fraser et al. (2024) — albeit using more sensitive data. A coarse summation presented here is sufficient for TESS due to its large pixel scale.
2.4 Recovery of Known Minor Planet 896 Sphinx
We validated our algorithm by recovering another small body which was observed in the same field as 3I/ATLAS. We find that the known main-belt minor planet 896 Sphinx (A918 PE) (H) was observed on Camera 2 CCD 3. 896 Sphinx was initially discovered in 1918 by M. Wolf. It has a known diameter of 11.9 km (J. R. Masiero et al., 2014) and rotation period of hours (T. Polakis, 2018). We apply our background-correction and shift-stack method outlined above to the region where 896 Sphinx is within the FFIs.
We present the deepstack image of 896 Sphinx in Figure 4. We find that we are able to recover 896 Sphinx at . Additionally, we extract a light curve using a -pixel mask centered on 896 Sphinx. We present the extracted normalized light curve and associated errors in Figure 4. We run this light curve through a Lomb-Scargle periodogram (N. R. Lomb, 1976; J. D. Scargle, 1982) to recover the known rotation period of the object. We find that periods hours are dominated by instrumental noise. We ignore these periods. We find the peak of max power for hours. We fit this peak with a Gaussian function and optimize our fit using a non-linear least squares approach. We find that the peak occurs at hours. Our error is adopted from the width of the best-fit Gaussian. Our recovered period is within of the period presented in T. Polakis (2018). The discrepancy between our rotation period and the archival rotation period may be due to the lack of sensitivity of TESS for faint objects.
2.5 Recovery of 3I/ATLAS
Figure 5 shows the background-removed deepstack images and associated errors of 3I/ATLAS as observed by TESS. These deepstacks are limited to cutouts which are not overcrowded with bright sources, as defined in Section 2.2. The median, mean, and summed images are shown for both detectors. The detection threshold for 3I/ATLAS is calculated by propagating errors from the calibrated FFIs. Although the median deepstack images are implemented for the remainder of this paper, it should be noted that 3I/ATLAS is visible by-eye in all three types of images in Figure 5. The central pixel in Camera 2 CCD 3 has a median counts value of e- s-1. The background scatter is calculated as the median and standard deviation of the singular pixels surrounding the central pixel exercise. The background has an averaged count value of e- s-1. Therefore, we are able to recover 3I/ATLAS at in this summed deepstack image. The same exercise is repeated for Camera 1 CCD 2. The central pixel in this image has a summed counts value of e- s-1. The background has an averaged count value of e- s-1. Therefore, we are able to recover 3I/ATLAS at in this summed deepstack image. It is not surprising that we are unable to recover 3I/ATLAS at the same high fidelity as the previous image. This is because the field becomes increasingly more crowded as 3I/ATLAS’ apparent motion places it closer to the galactic plane.
3 Results
3I/ATLAS has displayed visible cometary activity months prior to perihelion. It is feasible that the level of activity has changed and will continue to change throughout its orbit. Light curves from images when 3I/ATLAS was at further heliocentric distances could provide insights into variations in the activity and nucleus properties such as the rotation period. In this section, we therefore extract the 20-day TESS light curve and calculate the resulting secular light curve of 3I/ATLAS.
3.1 Magnitude Calculation
We use a pixel aperture on the median deepstacked image to calculate the magnitude of 3I/ATLAS. We convert the calibrated flux in e- s-1 to TESS magnitude using the following relationship:
In Equation 1, is the number of counts per second. This conversion is provided in the TESS Instrumental Handbook (R. Vanderspek et al., 2018). We adopt the error on the zeropoint of mag. We note that while the zeropoint may change slightly between sectors, cameras, and/or CCDs, it is standard practice to not recalculate the value for each observation combination (e.g. T. L. Farnham et al., 2019; M. M. Fausnaugh et al., 2023; J. E. Andrews et al., 2025). We calculate using a aperture on the median deepstack images presented in Figure 5. For Camera 2 CCD 3, we derive e- s-1. For Camera 1 CCD 2, we derive e- s-1.
We find that 3I/ATLAS has a in the Camera 2 CCD 3 image and a in the Camera 1 CCD 2 image. Our calculated is in agreement with J. Martinez-Palomera et al. (2025). Examining the origins of slight offsets between our extracted values is beyond the scope of this work. These two magnitudes are averaged over the two observing windows in Table 1, which span a range in heliocentric distance of 6.4 to 5.4 au, respectively. In that same time period the distance between 3I/ATLAS and the TESS spacecraft decreased by 0.9 au. The rapid change in both these properties can be used to constrain any changes to the overall activity of 3I/ATLAS between both times.
From geometric effects alone, and ignoring the small () phase angle change, we would expect the flux from 3I to increase by a factor of 1.5. Converting this back into magnitudes we expected 3I/ATLAS to have a measured magnitude of 20.5 in the Camera 1 CCD 2 images, rather than the observed of 19.28, indicating a factor of 5 increase in the flux, rather than 1.5. Within our uncertainty on the TESS magnitude, our observations are statistically inconsistent with an asteroid-like reflectance model that ignores dust production.
3.2 Photometric Light Curve
We extract a light curve using the same -pixel aperture and propagate the errors from the calibrated FFIs. We normalize the light curve. Additionally, we bin our light curve to 36-minute bins. We analyze the binned light curve with a Lomb-Scargle periodogram to search for evidence of periodicity. We do this for each light curve independently. We search for a rotation period between hours, which is mostly consistent with of other kilometer scale small bodies in the solar system (B. D. Warner et al., 2009). Both of the light curves and affiliated periodograms are presented in Figure 6.
There are several similarities between the Lomb-Scargle periodograms. First, we find that hours are dominated by noise and/or unremoved systematics. Second, we see that there is no agreement between periodograms at hours. Third, we see a some strong periodicity between 10-30 hours. This is more apparent in the light curve from Camera 1 CCD 2, when 3I/ATLAS 0.4 au closer. In particular, we find a peak in the periodogram close to the recently recovered period of 16.16-16.79 hours presented in R. de la Fuente Marcos et al. (2025); T. Santana-Ros et al. (2025).
To further assess this potential rotational signature, we run the same Lomb-Scargle test for a nearby background pixels in the FFI. Because the background can rapidly vary across the TESS detector, we select an example background pixel within -pixels of 3I/ATLAS. The results of this test are presented in Figure 7. We find that the background pixel periodogram for Camera 2 CCD 3 shows a similar peak at hours, but we do not see a similarly strong peak around this time for Camera 1 CCD 2. However, there are stronger periodic peaks in this periodogram, so it is challenging to assess whether or not this 16 hour period is physically motivated by the TESS observations alone.
We complete this exercise for all of the pixels within a region around 3I/ATLAS. We pay particular attention to pixels that lead and trail the source on sky, as they should reveal systematics produced by nearby stars. The leading and trailing pixels are defined as those diagonal to the center from the bottom left to the upper right (the direction of motion of 3I/ATLAS). The light curve extracted from each pixel is binned at a 36 minute resolution. A Lomb-Scargle periodogram is performed on each resulting light curve across the same frequency range (Figure 8). There are some similarities between the leading/trailing pixels that do not align with the periodic signals observed from 3I/ATLAS. In particular, the trailing/leading pixels display a similar periodic signal at 29 hours in the data from Camera 2 CCD 3. This signal is offset from a peak in the 3I/ATLAS observations by hours. There are similar offsets between the trailing/leading and 3I/ATLAS pixels for Camera 1 CCD 2.
Given the similarities in noise properties of the periodograms, it is challenging to draw any confident conclusions about 3I/ATLAS. Therefore, we conclude that these TESS light curves do no definitively capture a from the nucleus. This may be due to ongoing outgassing even at large heliocentric distances or due to the sensitivity of TESS, making any possible magnitude changes from the baseline difficult to definitively measure.
3.3 Secular Light Curve
We calculated the inferred absolute V magnitude of 3I/ATLAS from these observations to add to the light curve presented in D. Z. Seligman et al. (2025). The secular light curve of a comet provides insight into the type of volatile driving activity — or at least the volatility of the species driving activity. For example, M. Womack et al. (2021) presented a 4 year secular light curve of C/1995 O1 (Hale-Bopp) to infer that CO was responsible for the activity exterior to 2.6-3.0 au, while H2O was responsible for it interior to that.
We convert the calculated TESS magnitude to absolute V magnitude, , using the following. T. L. Farnham et al. (2021) derived a rough relationship between the apparent visual magnitude, , and TESS magnitude of
This relationship was derived using the Web TESS Viewing Tool assuming typical comet colors derived from the average of several Jupiter Family comets, long-period comets, and active centaurs. This relationship has an uncertainty of mag, which is dominated by the uncertainty in comet colors. We convert to using the following:
In Equation 3, is the comet-Sun distance, is the comet-TESS distance, and is an activity index. has a typical range from , where assumes the body is inactive and assumes the object has strong activity (E. Everhart, 1967). Given that 3I/ATLAS is at a distance of 6.4 au in these images, we assume .
We use the average and for each Camera/CCD configuration presented in Table 1. From our calculated TESS magnitude, following Equations 2 and 3, and adopting the uncertainty presented in T. L. Farnham et al. (2021), we derive and . We plot discovery and precovery ZTF observations of 3I/ATLAS in Figure 9.
4 Discussion
We present two deepstack images and a 20-day light curve of 3I/ATLAS as observed with NASA’s Transiting Exoplanet Survey Satellite. These data were obtained from 07 May to 03 June 2025, predating the initial discovery by nearly two months (L. Denneau et al., 2025). From our two epochs of these data we calculate that 3I/ATLAS had a and , the latter of which is consistent with more recently obtained observations from the TRAPPIST-North and TRAPPIST-South telescopes (D. Z. Seligman et al., 2025). Moreover, we identify a periodic signal at hours in the light curve from Camera 1 CCD 2. However, we find no statistically significant evidence of a rotation period from the nucleus when comparing the light curve of 3I/ATLAS to a neighboring background pixel. The lack of observed rotation may be due to the presence of a coma, uncertainties in the size of the nucleus if there is no coma, or decreased sensitivity of TESS to faint objects.
The derived from these data indicate that 3I/ATLAS brightened more than can be expected from the decrease in helio- and TESS-centric distances between the two periods. It is worth stating that the object traversed nearly 1 au during the course of these observations. We note that the absolute magnitude of 3I/ATLAS in our first epoch,=13.9, is consistent with activity from a small (R2.8 km, 15.4) nucleus derived from HST WFC3 imaging (D. Jewitt et al., 2025). This provides tentative evidence that 3I/ATLAS was active during the fortuitous TESS observations prior to its discovery on UT 01 July 2025. However, the 3 error technically permits the possibility that 3I/ATLAS was fainter by 1 magnitude during these observations. This is due to the uncertainties in the conversion from magnitudes to of 0.3 (R. Vanderspek et al., 2018; T. L. Farnham et al., 2021). Therefore, we conclude that these observations are consistent with the hypothesis that 3I/ATLAS was weakly active prior to its discovery. Future work could include deriving a new relationship between and other bandpasses for ISOs.
This distant activity of 3I/ATLAS — if confirmed — could be indicative of mass loss driven by a mechanism other than the sublimation of H2O ice. One possibility is that 3I/ATLAS is enriched in hypervolatiles such as CO or CO2. 2I/Borisov displayed a higher production of CO than H2O when these two quantities were measured contemporaneously (D. Bodewits et al., 2020; M. A. Cordiner et al., 2020; Z. Xing et al., 2020), although no direct measurement of CO2 has been successfully obtained in an interstellar object to date.333Production rates of CO2 have been inferred from the ration of O I emission features in the visible, A. McKay et al. 2024 It is also possible that this activity is driven by a more exotic hypervolatile such as H2 or N2. The nongravitational acceleration of 1I/‘Oumuamua energetically would have required the sublimation of an ice more volatile than H2O (Z. Sekanina, 2019; D. Seligman & G. Laughlin, 2020). It has therefore been proposed that this acceleration was driven by outgassing of N2 (A. P. Jackson & S. J. Desch, 2021), CO (D. Z. Seligman et al., 2021), or radiolytically produced H2 (J. B. Bergner & D. Z. Seligman, 2023). It is possible that distant activity of 3I/ATLAS is driven by one of these, but determining which neutral is responsible requires spectroscopic confirmation or long-term lightcurve modeling (E. Bufanda et al., 2023).
The change in absolute magnitude between our two deep stacks indicates a change in activity, likely the result of a rapidly warming thermal environment. Continued searches for precovery images and extensive monitoring of 3I/ATLAS prior to perihelion will be essential for characterizing this trend. In particular, other precovery images which may include stellar occultations of appulses (J. L. Ortiz et al., 2023; C. L. Pereira et al., 2025), especially from multi-filter observations, could constrain the coma density and composition. Moreover, spectroscopic follow-up observations could also confirm the existence of hypervolatiles in 3I/ATLAS.
Observations of future ISOs could also shed light on the extent to which this population displays distant activity and is enriched with hypervolatiles. Existing all-sky surveys, including the new Rubin Observatory Legacy Survey of Space and Time (LSST) have been predicted to detect future ISOs (A. Moro-Martín et al., 2009; T. Engelhardt et al., 2017; N. V. Cook et al., 2016; D. Seligman & G. Laughlin, 2018; D. J. Hoover et al., 2022; D. Miller et al., 2022; D. Landau et al., 2023; D. Marčeta, 2023; D. Marčeta & D. Z. Seligman, 2023; R. C. Dorsey et al., 2025). Reorienting TESS to observe the ecliptic for longer periods of time in future extended missions could aid in the precovery observations of newly detected ISOs.
5 Data and Software Availability
The calibrated TESS FFIs are available for bulk download on the Mikulski Archive for Space Telescopes.444https://archive.stsci.edu/tess/bulk_downloads/bulk_downloads_ffi-tp-lc-dv.html All of the analysis Python scripts, data behind the figures, and Python scripts for reproducing the figures presented here are available on GitHub.555https://github.com/afeinstein20/atlas-tess Additional, larger, data products will be made available on Zenodo upon publication.
6 Acknowledgments
We thank Henry Hsieh, Davide Farnocchia, and Marco Micheli for thoughtful insights. We thank the anonymous referee for their comments which have improved the quality and clarity of this manuscript. A.D.F. acknowledges funding from NASA through the NASA Hubble Fellowship grant HST-HF2-51530.001-A awarded by STScI. D.Z.S. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303553. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time.
7 Author Contributions
A.D.F. lead the TESS data reduction, the creation of the deepstack images, the light curve extraction, manuscript writing, and figure creation. J.W.N. contributed manuscript text and review. D.Z.S. contributed manuscript text and review.
Appendix A Background Removal
In Figure 10 we present the models and residuals tested to remove the background properties in the TESS FFIs. These experiments demonstrated that the Savitsky-Golay filter (with a window length of 307) presents the optimal fit to the observations without overfitting the faint signal of 3I/ATLAS for Camera 2 CCD 3. The larger window length is also able to recover the target but improperly fits the sharp decline of the orbital ramp. The 2nd-order polynomial does not fit the Camera 1 CCD 2 data well due to the strong systematics present in the observations. Similarly to Camera 2 CCD 3, the second image is best fit by the Savitsky-Golay filter with a window length of 307. Therefore, we select this background to subtract from the calibrated FFIs in the entirety of the analysis presented in this work.
Appendix B Aperture Selection
The TESS pixel response function (PRF) can change substantially over a given camera’s FOV. Additionally, the PRF is slightly chromatic and can vary with temperature. The TESS PRF model assumes that 24% of the total source flux will be contained in the central pixel if the target is perfectly centered. If the target is offset, then the total amount of flux decreases. It is particularly important to take this into account when trying to extract the flux from a moving target. To combat these effects, larger apertures are more robust against these flux losses. However, implementing such large apertures is also particularly challenging for crowded fields because they may include flux contributions from nearby sources.
To investigate the efficacy of both competing effects, we perform a set of experiments on a variety of apertures to extract the light curve of 3I/ATLAS. The subset of apertures investigated here are provided in A. D. Feinstein et al. (2019), ranging from a single pixel to a area. The results of these tests are presented in Figure 11 and Figure 12. These figures show the full light curve as well as the light curve binned to 36 minutes. We also run the binned light curve through a Lomb-Scargle periodogram to search for potential evidence of a rotation period.
Figure 11 demonstrates that the light curve is dominated by noise at hours. Additionally, there is no strong evidence of periodicity for hours. While there are stronger peaks at hours, the majority of these periodograms have peaks at identical times within this time range. We note that the periodograms shown in rows 5-8 of Figure 11 contain similar peaks at hours, which could be consistent with the reported hours (R. de la Fuente Marcos et al., 2025). However, because the same signal is not seen across all extracted light curves, we attribute it to instrumental systematics.
We repeat this same test with the data extracted from Camera 1 CCD 2. These results are presented in Figure 12. Similarly to the previous analysis of the other TESS image, the hours is dominated by noise. Contrary to the previous analysis however, there is a strong peak between hours. Additionally, we find a consistent strong double-peaked feature at hours. The 16 hour peak is consistent with recovered by (R. de la Fuente Marcos et al., 2025; T. Santana-Ros et al., 2025). These signals are strong in every extracted light curve. None of the stronger peaks seen in the light curves extracted from Camera 1 CCD 2 are consistent with those from Camera 2 CCD 3. Additionally, because there are multiple strong peaks and no observable photometric variability in all data, it is challenging to determine whether the signals are instrumental or astrophysical in origin. Nevertheless, because the periodograms between TESS data sets are inconsistent, we conclude that we are unable to measure a rotation period from these observations.